OpenStudy (anonymous):

Medal to the best answer! A drawer contains 4 red socks, 3 white socks, and 3 blue socks. Without looking, you select a sock at random, replace it, and select a second sock at random. What is the probability that the first sock is blue and the second sock is red?

jimthompson5910 (jim_thompson5910):

what's the probability of picking blue

OpenStudy (anonymous):

3/10?

jimthompson5910 (jim_thompson5910):

there are 3 blue socks there are 4+3+3 = 10 socks total so the probability of picking blue is 3/10 good catch

jimthompson5910 (jim_thompson5910):

what's the probability of picking red

OpenStudy (anonymous):

4/10

OpenStudy (anonymous):

would the answer be 7/20?

jimthompson5910 (jim_thompson5910):

so from here, you multiply the two probabilities because the two events are independent P(Blue then red) = P(Blue AND Red) P(Blue then red) = P(Blue)*P(Red) P(Blue then red) = (3/10)*(4/10) P(Blue then red) = (3*4)/(10*10) P(Blue then red) = 12/100 P(Blue then red) = 3/25

OpenStudy (anonymous):

Thanks @jim_thompson5910 !

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (kenljw):

by multiplication rule there are 4X3X3 outcomes P(R) = 4!/4X3X3=4X3X2/4X3X3=2/3 P(B) = 3!/4X3X3 = 3X2/4X3X3 = 1/6 use 3! because there 3X2X1 ways of selecting them conditional probability P(R|B) = P(R) because you used replacement P(B and then R) = P(B)P(R|B) = 2/9

OpenStudy (kenljw):

2/18